# Literature Review: Cyber Security Monitoring in Maritime

### v1

@@ V-D Security monitoring methods for maritime CPSs
Pelissero, Laso and Puentes [44] created and utilised graph-based models for the detection of attack propagation in CPSs. A graph model that represented the digital and physical system layers was proposed in the paper. It considered the interdependencies between CPSs by using the system variables. The model was applied in a case study that involved a small-scale system consisting of a ship with a fuel tank, valve, motor and propeller which were monitored through a PLC. One scenario of normal system usage and four attack scenarios including port scan, closing the PLC and man-in-the-middle attacks were simulated in the study. According to the authors, the method did not produce a false positive alert for the normal scenario, whereas one attack scenario out of four remained undetected (i.e., precision and recall were 100% and 75%, respectively). However, the evaluation of the method was limited to aforementioned five scenarios without testing the method on a larger dataset. The use of a significantly larger dataset would provide a better estimate about the long-term performance of the method when cyber attacks of a number of different types need to be handled. Furthermore, the applicability of the method would be questionable in large-scale systems as identification of the dependencies between system components is a non-trivial task that involves dedicated approaches [56].
@@
Yoginath et al. [46] proposed a Recurrent Neural Network (RNN) based Digital Twin (DT) model to simulate an operational CPS of a Canal Lock. Their model performed real-time analysis to detect anomalies at the earliest, specifically anomalous changes in water levels within chambers through which ships pass. The study [46] used an actual PLC, sensors, valves and pumps to emulate the behaviour of the Canal Lock CPS physically. The authors identified all the possible scenarios, i.e., the ship’s up and down movements, performed by Canal Lock CPS and trained a set of RNN models using the emulation data as input. Anomaly detection was performed through a difference-based method (analysis of PLC and DT predicted values). The previously DT predicted value was compared with the current PLC read value and the deviation was captured, with the larger value being regarded as an indication of the attack. Furthermore, to evaluate the RNN model, the authors performed live validation of the system. However, any specific CPS attacks were not executed to test the accuracy of the DT concept for anomaly detection.
@@
Xing, Cao and Chen [51] presented an anomaly detection method for the Ship Information System (SIS) based on risk data analysis. Since SIS is a typical SCADA system, the authors first elaborated on all the different operational and networking components of the SIS SCADA system and then they proceeded with presenting the data anomaly detection approach based on a cooperative state space control mathematical model, for the data propagated between the sensors, the distributed controller units and the actuators. A so-called Critical State Estimation (CSE) Algorithm was proposed based on the Industrial State Modelling Language (ISML). Detection rules were formulated as “condition implies action” with the condition being a boolean formula composed by various predicates indicating values that are expected by system components. The approach was validated by simulations on the ship dynamics under heading sensor signal attacks. The implementation of the controller was not made publicly available nor was the validation sensor data.
@@
The study by Boudehenn et al. [43] focused on the use of Teager-Kaiser operator for time series analysis to detect anomalies in a ship’s propulsion system in the case of cyber attacks. During the experiments, the authors collected the following time series data – oil consumption, fuel consumption, propeller speed, and valve opening. The following attacks were conducted for evaluating the anomaly detection methods – port scanning with nmap tool, DoS attack with hping tool, and exploiting the vulnerability of the ship’s propulsion PLC for stopping and starting the PLC. Port scanning and DoS attacks did not influence the propulsion system, and only the attack against the PLC had a significant impact. According to the experiments, the Teager-Kaiser operator managed to identify sudden changes in the ship’s propulsion system time series data. However, the evaluation results were presented in the form of graphs only for the 5-minute time period which involves a single attack against the PLC. Therefore, it is difficult to assess what would be the detection rate for many attacks of different types against the propulsion system during a longer time frame (e.g., several days or weeks). Furthermore, the study did not include the analysis of false positive rates over a longer period of normal navigation without any cyber attacks (normal navigation can easily involve sudden changes in the propulsion system time series data, e.g., due to changes in engine speed by the crew). Although the study used a maritime specific dataset, neither the data nor the experiment code were released to the public domain by the authors. In addition, the study did not present any data on the computational cost of the anomaly detection method.
@@
The study by Qiu et al. [41] proposed the use of a Particle filter-based anomaly detection method for time series data collected from a ship’s CPS to detect cyber attacks. In the study, the authors focused on data collected from the ship’s propulsion system – the temperature of the cylinder exhaust, cooling water pressure, and the speed of the fuel flow. For training the anomaly detection system, the authors used data collected from a new cargo ship during its first sailing week, assuming that the probability of cyber attacks during this week is very low. For evaluating the anomaly detection algorithm, data from the second half of the first sailing year was employed. For discussing the anomaly detection rate of the algorithm, the authors presented graphs for the 5,000-minute time frame (about 3.5 days) in the study. Although the graphs displayed some fluctuations in the time series data for the propulsion system, the study did not detail whether these sudden changes were the result of a cyber attack or rather a side effect of normal navigation. Also, the paper did not provide any experimental data that would involve real or simulated cyber attacks, measuring the precision and recall of the proposed anomaly detection method (traditional metrics for characterizing the performance of anomaly detection methods). In addition, the study did not provide any information about the computational cost of the method. Also, neither the experiment data nor the implementation of the anomaly detection method were released to the public domain.
@@
As for the types of security monitoring methods (see Table IV), no particular method types have been dominantly employed for security monitoring of maritime CPSs, and proposed methods have included supervised ML [41, 46], unsupervised ML [43], rule-based [51], and graph-based [44] methods. However, when analysing OT data from maritime CPSs, existing methods have generally not considered discriminating cyber attack related CPS malfunctions from CPS faults not caused by malicious cyber activity. Therefore, the root cause analysis for maritime CPS faults and anomalies is an unexplored research area which deserves closer attention.
@@
Finally, in order to properly study the effects of cyber attacks on maritime CPS, a highly realistic lab environment is needed which would mimic a real-life CPS as closely as possible. Apart from one recent work [39] that will be discussed in Section V-F, the creation of realistic maritime CPS testbeds has been understudied, and can be regarded as a promising future research direction.
@@ V-E Security monitoring system architectures
The proposed shore-based functional blocks of the SOC were the following. First, Ship Shore Manager and Bandwidth Manager were responsible for receiving the security data from ships and managing network connectivity with remote parties; Central Processor was responsible for filtering and normalising the received security data, and Data Store was a big data solution that stored the normalised security data. In addition to the aforementioned blocks, the shore-based SOC had other functional blocks that can also be found in traditional SOCs. For example, human operators could use specialised tools for big data analytics, searching the collected security data, creating visualizations, etc., to achieve maritime cyber situational awareness. In addition, SOC could be linked to external threat intelligence sources and incident response platforms. Finally, collected security data could be used for creating a so-called maritime-recognised cyber picture of the status of individual ships and the entire fleet. In their study [13], the authors mentioned that the previously described architecture has been thoroughly tested. Although no detailed information was provided about the nature of these tests and the performance of the architecture, the architecture was validated in a later independent study [14] by building a maritime exercise environment according to this architecture (see Section V-F).
Schauer et al. [47] proposed a security monitoring system architecture for ports which would integrate security data from two distinct domains – traditional IT systems (i.e., cyber assets) and physical assets of the port. The authors proposed to collect a wide variety of security events from IT systems, including firewall events, antivirus events from end user devices, and event logs of applications and servers. To monitor the physical assets of the port, the authors suggested data collection from physical access control systems, dedicated monitoring sensors which can create security events, etc. According to the authors, security events from cyber and physical assets should be collected into the hybrid situational awareness framework which can analyse events from both domains.
@@
For real-time analysis of these events, two major components were proposed for the situational awareness framework – Event Correlation Engine (ECE) and Threat Propagation Engine (TPE). According to a widely used definition [58], event correlation is a real-time process that involves assigning a new meaning to event groups which occur within predefined time windows. For implementing ECE, the DROOLS engine was proposed in [47]. As the authors stated, the use of ECE allows the detection of complex security incidents which manifest themselves through several events from different domains. The purpose of the TPE was to establish the impact of high-priority security alarms via graph-based analysis, where the nodes of the graph represented cyber and physical assets and the edges of the graph reflected the dependencies between assets.
The approach by Schauer et al. [47] was sound and novel considering the time of its publication, integrating cyber and physical situational awareness methods for port infrastructures (including communication with ships). On the other hand, the authors did not provide a deeper analysis of the possible cyber attacks and how they could be detected in the maritime infrastructure domain. Also, the approach was heavily human expert-based with regards to the creation of the pattern-based rules for the event correlation engine and also to raising alarms for triggering the operation of the threat propagation engine. This may contribute to biased decision making. Furthermore, the authors proposed the DROOLS engine for event correlation purposes without comparing it with other solutions (e.g., [59, 60]) to justify the selection of DROOLS. Last but not least, the approach was exemplified through two attack scenarios but no simulation results were provided and the starting of the attacks was assumed to succeed rather easily.
@@
The paper by Jia et al. [50] discussed a big data architecture to collect OT data from Unmanned Surface Vehicles (USVs) for various purposes like business management, disaster prevention, anomaly detection, etc. The authors proposed the collection of the following data from USVs – ship’s performance and navigational data (e.g., speed, position, level of fuel, etc.), data about the environment (e.g., wind, temperature, etc.), data from sensing devices that provided information about the objects around the ship (e.g., data from the radar), and data that was collected from other sources than USV itself (e.g., industrial data about the USV collected from the manufacturer). For storing the collected data, a cloud-based solution was proposed that was PostgreSQL based. According to the authors, visualisation capabilities were an important aspect of the big data architecture, and the authors suggested the use of the LinDA toolkit for processing the data in the PostgreSQL database.
As an example use case for the big data architecture, Jia et al. described an anomaly detection scenario for identifying abnormal combinations of the sailing angle and the USV speed. For anomaly detection purposes, Isolation Forest was proposed. However, the authors did not provide a performance evaluation of the Isolation Forest based anomaly detection method on some datasets. Also, the code and the data used during the experiments were not made publicly available by the authors.
@@
In [48] Möller et al. discussed an agent-based intelligent maritime traffic management system to support a coastal Vessel Traffic Service (VTS) with a clear situational picture, traffic and environmental information. For that purpose, the authors proposed to introduce a Multi-Agent System (MAS) that extended to vessels to exchange relevant data (e.g., navigational and safety information) with the central VTS. Following the concept of connected cars, the authors also introduced the concept of connected ships which were continuously in contact with each other. That facilitated the exchange of important navigational information not only through VTS but directly between ships.
The authors also described how to augment the proposed distributed MAS with a security monitoring functionality. First, the authors detailed the weaknesses of the wireless communication networks used by MAS, and mentioned a mitigation opportunity in the form of anomaly detection for data flows in the network. However, the authors did not provide a more detailed discussion on the anomaly detection algorithm suitable for that goal. Möller et al. also suggested the use of neural networks for the detection of cyber attacks. However, the discussion of the proposed method remained brief and no implementation information was provided. For example, although the study mentioned training neural networks on probability-based features, the nature of the input data was not detailed. Furthermore, the paper did not report any experiments with the proposed method together with attack detection performance results. For these reasons, it is difficult to assess if the suggested neural network based cyber attack detection method would be a viable choice for maritime systems.
@@
According to recent studies [62, 16], SOC analysts are often overwhelmed by security alerts of low importance which leads to analyst burnout and alert fatigue. For supporting the human analysts, alert prioritisation has been identified as an important topic for SOC environments [16], and several ML approaches have been recently proposed for that purpose [63, 64]. Since the employment of alert prioritisation for reducing human alert fatigue has not been studied in the context of maritime SOC, it can be regarded as one of the open issues for future work.
@@ V-F Security monitoring testbeds
The short paper authored by Becmeur et al. [52] aimed to present a platform to generate data and scenario traces for the evaluation of algorithms for intrusion detection. The introduced solution was based on a ship’s two critical subsystems: 1) The propulsion and engine control in which a computer fan simulated the ship’s propeller and the fan’s actual RPM represented propulsion. 2) The navigation subsystem with the rudder control to simulate the ship’s direction changes. Each subsystem had an independent controller which used common industrial communication protocols, like Modbus, DNP3, and S7. The heart of the setup was the central controller built on a Raspberry Pi. The simulated sensor – that provided the measurement of the engine parts’ temperature – was made of an open-source electronics platform called Arduino. The authors provided an overview of the project, elucidating its overarching objectives and outcomes, but the overview lacked the required depth in elucidating the procedural intricacies that are essential for comprehensively understanding and, crucially, reproducing the results.
@@
The paper by Sicard et al. [39] described a testbed for experimenting with the Industrial Control System (ICS) of a warship. To achieve a high degree of realism, the authors utilised the physical devices of real ships as much as possible. The testbed implemented the following four areas of the warship – propulsion system (engine and propellers), direction (rudders), energy (fuel for the engine and fuel pump), and artillery (76mm main gun and motors for moving the gun turret). For these four areas, a physical implementation contained a bow and rear of the ship with a physical gun turret, rudders, propellers, physical devices of the bridge, etc. According to the authors, the use of virtualised solutions would have decreased the realism of the testbed, since they would have not allowed for the full imitation of the ship’s real-life environment. For low-level monitoring of Modbus RTU fieldbus traffic in the ICS of the ship, the authors proposed the use of Zeek IDS with specialised traffic capturing hardware and Zeek protocol parser for the Fieldbus protocol. For regular network monitoring, commercial IDS sensors were suggested by the authors. Although the study mentioned that the testbed could be utilised for generating maritime cyber security datasets and described four possible cyber attack scenarios, such datasets were not published by the study.
@@
Discussion. From the three analysed papers, two papers have proposed testbeds for experimenting with vessel ICS and SCADA networks [52, 39], whereas [14] focuses on a SOC testbed for training maritime SOC analysts. Since [52] is a short paper, the discussion of the proposed testbed remains fairly brief, but in the remaining two papers [14, 39] detailed testbed descriptions are provided, with both testbeds being highly realistic environments for conducting maritime cyber security experiments and trainings.
@@
However, it should be considered that the aforementioned studies represent specific environments (a warship’s ICS and a maritime SOC). Since for other maritime systems there are no similar studies, the creation of highly realistic testbeds for these systems constitutes an open research area. Also, realistic testbeds allow for generating high-quality datasets for maritime cyber security experiments. As discussed in Section V-A, there are very few publicly available maritime-specific datasets (see Table V), and the creation of such datasets is another important future research direction.

### v2

@@ V-D Security monitoring methods for maritime CPS
Pelissero et al. [49] created and utilised graph-based models for the detection of attack propagation in CPSs. A graph model that represented the digital and physical system layers was proposed in the paper, considering the interdependencies between CPSs by using the system variables. The model was applied in a case study that involved a small-scale system consisting of a ship with a fuel tank, valve, motor and propeller which were monitored through a PLC. One scenario of normal system usage and four attack scenarios including port scan, closing the PLC and man-in-the-middle attacks were simulated in the study. According to the authors, the method did not produce a false positive alert for the normal scenario, whereas one attack scenario out of four remained undetected (i.e., precision and recall were 100% and 75%, respectively). However, the evaluation of the method was limited to the aforementioned five scenarios without testing the method on a larger dataset.
@@
Yoginath et al. [51] proposed a Recurrent Neural Network (RNN) based Digital Twin (DT) model to simulate an operational CPS of a Canal Lock. Their model performed real-time analysis to detect anomalies at the earliest, specifically anomalous changes in water levels within chambers through which ships pass. The study [51] used an actual PLC, sensors, valves and pumps to emulate the behaviour of the Canal Lock CPS physically. The authors identified all the possible scenarios, i.e., the ship’s up and down movements, performed by Canal Lock CPS and trained a set of RNN models using the emulation data as input. Anomaly detection was performed through a difference-based method (analysis of PLC and DT predicted values). The previously DT predicted value was compared with the current PLC read value and the deviation was captured, with the larger value being regarded as an indication of the attack. Furthermore, to evaluate the RNN model, the authors performed live validation of the system. However, no specific CPS attacks were executed to test the accuracy of the DT concept for anomaly detection.
@@
Xing et al. [56] presented an anomaly detection method for the Ship Information System (SIS) based on risk data analysis. Since SIS is a typical SCADA system, the authors first elaborated on all the different operational and networking components of the SIS SCADA system and then they proceeded with presenting the data anomaly detection approach based on a cooperative state space control mathematical model, for the data propagated between the sensors, the distributed controller units and the actuators. A so-called Critical State Estimation (CSE) Algorithm was proposed based on the Industrial State Modelling Language (ISML). Detection rules were formulated as “condition implies action” with the condition being a boolean formula composed by various predicates indicating values that are expected by system components. The approach was validated by simulations on the ship dynamics under heading sensor signal attacks.
@@
The study by Boudehenn et al. [48] focused on the use of Teager-Kaiser operator for time series analysis to detect anomalies in a ship’s propulsion system in the case of cyber attacks. During the experiments, the authors collected the following time series data – oil consumption, fuel consumption, propeller speed, and valve opening. The following attacks were conducted for evaluating the anomaly detection methods – port scanning with nmap tool, DoS attack with hping tool, and exploiting the vulnerability of the ship’s propulsion PLC for stopping and starting the PLC. Port scanning and DoS attacks did not influence the propulsion system, and only the attack against the PLC had a significant impact. According to the experiments, the Teager-Kaiser operator managed to identify sudden changes in the ship’s propulsion system time series data. However, the evaluation results were presented in the form of graphs only for the 5-minute time period which involved a single attack against the PLC. Therefore, it is difficult to assess what would be the detection rate for many attacks of different types against the propulsion system during a longer time frame (e.g., several days or weeks), and a more detailed analysis of the proposed method was identified as future work in the study.
@@
The study by Qiu et al. [46] proposed the use of a Particle filter-based anomaly detection method for time series data collected from a ship’s CPS to detect cyber attacks. In the study, the authors focused on data collected from the ship’s propulsion system – the temperature of the cylinder exhaust, cooling water pressure, and the speed of the fuel flow. For training the anomaly detection system, the authors used data collected from a new cargo ship during its first sailing week, assuming that the probability of cyber attacks during this week is very low. For evaluating the anomaly detection algorithm, data from the second half of the first sailing year was employed. For discussing the anomaly detection rate of the algorithm, the authors presented graphs for the 5,000-minute time frame (about 3.5 days) in the study. Although the graphs displayed some fluctuations in the time series data for the propulsion system, the study did not detail whether these sudden changes were the result of a cyber attack or rather a side effect of normal navigation.
@@
As for the types of security monitoring methods (see Table III), no particular method types have been dominantly employed for security monitoring of maritime CPSs, and proposed methods have included supervised ML [46, 51], unsupervised ML [48], rule-based [56], and graph-based [49] methods. However, when analysing OT data from maritime CPSs, existing methods have generally not considered discriminating cyber attack related CPS malfunctions from CPS faults not caused by malicious cyber activity. Therefore, the root cause analysis for maritime CPS faults and anomalies is an unexplored research area which deserves closer attention.
@@
Finally, in order to properly study the effects of cyber attacks on maritime CPS, a highly realistic lab environment is needed which would mimic a real-life CPS as closely as possible. Apart from a few recent works [44, 38] that will be discussed in Section V-F, the creation of realistic maritime CPS testbeds has been understudied, and can be regarded as a promising future research direction.
@@ V-E Security monitoring system architectures
The proposed shore-based functional blocks of the SOC were the following. First, Ship Shore Manager and Bandwidth Manager were responsible for receiving the security data from ships and managing network connectivity with remote parties; Central Processor was responsible for filtering and normalising the received security data, and Data Store was a big data solution that stored the normalised security data. In addition to the aforementioned blocks, the shore-based SOC had other functional blocks that can also be found in traditional SOCs. For example, human operators could use specialised tools for big data analytics, searching the collected security data, creating visualisations, etc., to achieve maritime cyber situational awareness. In addition, SOC could be linked to external threat intelligence sources and incident response platforms. Finally, collected security data could be used for creating a so-called maritime-recognised cyber picture of the status of individual ships and the entire fleet. In their study [13], the authors mentioned that the previously described architecture has been thoroughly tested. Although no detailed information was provided about the nature of these tests and the performance of the architecture, the architecture was validated in a later independent study [14] by building a maritime exercise environment according to this architecture (see Section V-F).
@@
Schauer et al. [52] proposed a security monitoring system architecture for ports which would integrate security data from two distinct domains – traditional IT systems (i.e., cyber assets) and physical assets of the port. The authors proposed to collect a wide variety of security events from IT systems, including firewall events, antivirus events from end user devices, and event logs of applications and servers. To monitor the physical assets of the port, the authors suggested data collection from physical access control systems, dedicated monitoring sensors which can create security events, etc. According to the authors, security events from cyber and physical assets should be collected into the hybrid situational awareness framework which can analyse events from both domains. For real-time analysis of these events, two major components were proposed for the situational awareness framework – Event Correlation Engine (ECE) and Threat Propagation Engine (TPE). According to a widely used definition [64], event correlation is a real-time process that involves assigning a new meaning to event groups which occur within predefined time windows. For implementing ECE, the DROOLS engine was used. As the authors stated, the use of ECE allows the detection of complex security incidents which manifest themselves through several events from different domains. The purpose of the TPE was to establish the impact of high-priority security alarms via graph-based analysis, where the nodes of the graph represented cyber and physical assets and the edges of the graph reflected the dependencies between assets. The approach by Schauer et al. [52] was sound and novel considering the time of its publication, integrating cyber and physical situational awareness methods for port infrastructures (including communication with ships). On the other hand, the study did not provide a deeper analysis of the possible cyber attacks and how they could be detected in the maritime infrastructure domain, exemplifying the proposed approach through two attack scenarios.
@@
The paper by Jia et al. [55] discussed a big data architecture to collect OT data from Unmanned Surface Vehicles (USVs) for various purposes like business management, disaster prevention, anomaly detection, etc. The authors proposed the collection of the following data from USVs – ship’s performance and navigational data (e.g., speed, position, level of fuel, etc.), data about the environment (e.g., wind, temperature, etc.), data from sensing devices that provided information about the objects around the ship (e.g., data from the radar), and data that was collected from other sources than the USV itself (e.g., industrial data about the USV collected from the manufacturer). For storing the collected data, a cloud-based solution was proposed that was PostgreSQL based. According to the authors, visualisation capabilities were an important aspect of the big data architecture, and the authors suggested the use of the LinDA toolkit for processing the data in the PostgreSQL database. As an example use case for the big data architecture, Jia et al. described an anomaly detection scenario for identifying abnormal combinations of the sailing angle and the USV speed. For anomaly detection purposes, an Isolation Forest-based method was proposed, but the study did not include its performance evaluation.
@@
Möller et al. [53] discussed an agent-based intelligent maritime traffic management system to support a coastal Vessel Traffic Service (VTS) with a clear situational picture, traffic and environmental information. For that purpose, the authors proposed to introduce a Multi-Agent System (MAS) that extended to vessels to exchange relevant data (e.g., navigational and safety information) with the central VTS. Following the concept of connected cars, the authors also introduced the concept of connected ships which were continuously in contact with each other. That facilitated the exchange of important navigational information not only through VTS but directly between ships. The authors also described how to augment the proposed distributed MAS with a security monitoring functionality. First, the authors detailed the weaknesses of the wireless communication networks used by MAS, and mentioned a mitigation opportunity in the form of anomaly detection for data flows in the network. The study also suggested the use of neural networks for the detection of cyber attacks, but did not report any experiments with the proposed method together with attack detection performance results.
@@
According to recent studies [66, 16], SOC analysts are often overwhelmed by security alerts of low importance which leads to analyst burnout and alert fatigue. For supporting the human analysts, alert prioritisation has been identified as an important topic for SOC environments [16], and several ML approaches have been recently proposed for that purpose [67, 68]. Since alert prioritisation algorithms for reducing human alert fatigue have not been studied in the context of maritime SOC, it can be regarded as one of the open issues for future work.
@@ V-F Security monitoring testbeds
The short paper authored by Becmeur et al. [57] aimed to present a platform to generate data and scenario traces for the evaluation of algorithms for intrusion detection. The introduced solution was based on a ship’s two critical subsystems: 1) the propulsion and engine control in which a computer fan simulated the ship’s propeller and the fan’s actual RPM represented propulsion, 2) the navigation subsystem with the rudder control to simulate the ship’s direction changes. Each subsystem had an independent controller which used common industrial communication protocols, like Modbus, DNP3, and S7. The heart of the setup was the central controller built on a Raspberry Pi. The simulated sensor – that provided the measurement of the engine parts’ temperature – was made of an open-source electronics platform called Arduino. The authors provided an overview of the project, describing its overarching objectives and outcomes, but the overview lacked the required depth for allowing to recreate a similar platform.
@@
The paper by Sicard et al. [44] described a testbed for experimenting with the Industrial Control System (ICS) of a warship. To achieve a high degree of realism, the authors utilised the physical devices of real ships as much as possible. The testbed implemented the following four areas of the warship – propulsion system (engine and propellers), direction (rudders), energy (fuel for the engine and fuel pump), and artillery (76mm main gun and motors for moving the gun turret). For these four areas, a physical implementation contained a bow and rear of the ship with a physical gun turret, rudders, propellers, physical devices of the bridge, etc. According to the authors, the use of virtualised solutions would have decreased the realism of the testbed, since they would have not allowed for the full imitation of the ship’s real-life environment. For low-level monitoring of Modbus RTU fieldbus traffic in the ICS of the ship, the authors proposed the use of Zeek IDS with specialised traffic capturing hardware and Zeek protocol parser for the Fieldbus protocol (in [69], the authors provided a more detailed description of this solution). For regular network monitoring, commercial IDS sensors were suggested by the authors.
Basels et al. [36] introduced a testbed to identify navigation radar vulnerabilities and to experiment with relevant security monitoring solutions. The offensive module of the testbed was based on the Radar Attack Tool (RAT) [59], which is able to launch a variety of image manipulation attacks against maritime radars (e.g., freezing or scaling the image, removing an object from the image, etc.). The defensive module of the testbed involved Snort3 and IPAL IDS solutions which were configured to detect the attacks against maritime radars by analysing application layer data in network traffic. For Snort3, the authors developed the rules for detecting message injection attacks, whereas IPAL was employed for detecting message modification scenarios. The detection by IPAL relied on two methods outlined in [70].
The paper by Longo et al. [38] described the MaCySTe testbed designed for maritime cyber security experiments with the vessel IT/OT systems. The authors released detailed installation instructions of the testbed together with the source code in order to facilitate maritime cyber security research in other organisations. The testbed was built on top of Linux containers, making it a highly resource-efficient solution with modest computational requirements. The testbed provided a detailed emulation of the vessel network, navigational devices, propulsion system, and other components together with relevant protocols (e.g., Navico BR24 and Modbus). For evaluating security monitoring functionality, the testbed supported the use of network probes and SIEM. The network probes captured and parsed traffic in the vessel network, forwarding relevant data to SIEM for visualisation and analysis. The authors also provided some illustrative examples on how the MaCySTe testbed can be utilised for cyber attack and attack detection experiments.
@@
Discussion. From the five analysed papers, two papers have proposed testbeds for experimenting with vessel ICS and SCADA networks [57, 44], one paper has proposed a testbed for maritime radars [36], one paper has described a more general maritime testbed for a wide range of cyber security experiments [38], whereas [14] focuses on a SOC testbed for training maritime SOC analysts. Since [57] is a short paper, the discussion of the proposed testbed remains fairly brief, but in the remaining four papers [14, 44, 36, 38] detailed testbed descriptions are provided, with all testbeds being highly realistic environments for conducting maritime cyber security monitoring experiments and trainings.
@@
However, it should be considered that some of the aforementioned studies represent specific environments (e.g., warship’s ICS). Therefore, the creation of highly realistic maritime cyber security testbeds remains an open research area. Also, realistic testbeds allow for generating high-quality datasets for maritime cyber security monitoring experiments. As discussed in Section V-A, although recent studies have publicly released several maritime-specific datasets, their number remains limited (see Table IV), and the creation of such datasets is another important future research direction.